OpenStudy (anonymous):

will give fan Maggie babysat for 61/3 hours on Saturday night. She earns $15 per hour for babysitting. She plans to spend $7 of the money she earned to see a movie. Which number sentence could be used to find E, the amount of her earnings which will be left if the only thing she buys is the movie ticket? E = 61/3 × ($15 - $7) E = (61/3 × $15) - $7 E = ($15 ÷ 61/3) - $7 E = 61/3 + $15 - $7

OpenStudy (abbles):

E = 61/3 x 15 - 7 The third option.

OpenStudy (anonymous):

explain pleas

OpenStudy (abbles):

So Maggie earned 15 per hour babysitting, correct? To find how many hours she worked total, you would need to divide 61 by 3. That gives you the number of hours she worked which you would then need to multiply by $15 to get the total amount of money she earned. Then, she saw a movie and used 7 dollars which we would need to subtract from her total amount of money in order to see how much she has left. Do you understand?

OpenStudy (anonymous):

u were rong :'(

OpenStudy (abbles):

How?

OpenStudy (abbles):

Is this the one you chose? E = (61/3 × $15) - $7

OpenStudy (anonymous):

idk First, determine the total amount she earns from babysitting. The amount she earns is equal to the number of hours she works multiplied by the amount she charges per hour. 61/3 × $15 The amount of her earnings which she will have left is the cost of the movie ticket subtracted from her total earnings. (61/3 × $15) - $7

OpenStudy (abbles):

Right, so that should be correct, no?

OpenStudy (anonymous):

yes but it sead it was I guess it had a glitch I b your fan though

OpenStudy (abbles):

Do you have a teacher monitoring whatever program you're using? Maybe talk to them about it? I'm positive E = (61/3 × $15) - $7 is the correct answer, maybe make sure that's the one you chose. Thanks for the fan :) sorry it marked it wrong. Let me know if you have any other questions I could help out with.